Circular No. 6497

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bo IN NGC 673 A supernova in NGC 673 has been discovered independently by Wei-

dong Li, Qiran Qiao, Yulei Qiu, and Jingyao Hu, Beijing Astronomical Observatory (BAO), as part of the BAO Supernova Survey, on CCD images taken on Oct. 18; and by Mark Armstrong, Rolvenden, Kent, U.K. (communicated by G. M. Hurst, Basingstoke), on unfiltered CCD images taken on Oct. 23.095 UT. G. J. Garradd, Loomberah, N.S.W., reports that SN 1996bo is located ∼ 6′′ east of the galaxy’s nucleus at the following CCD position: α = 1^{h}48^{m}22^{s}.80, δ = +11^{o}31′15′′.8 (equinox 2000.0). The BAO observers note that their discovery was made on a 120-s exposure with the 0.60-m reflector, while a 300-s exposure on Oct. 8 with the same instrumentation showed no image at the position of SN 1996bo. Armstrong notes that images taken on Sept. 11 and Oct. 12 also failed to show anything at this position. Estimated magnitudes for SN 1996bo: Oct. 18, V ∼ 16.5 (Li et al.); 23.468, R ∼ 16.5 (Garradd).

M. Turatto and S. Benetti, European Southern Observatory (ESO); and M. Lemke, Sternwarte Bamberg, write: “A spectrum (range 360–740 nm, resolution 0.5 nm) taken on Oct. 24.10 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) confirms that this is a supernova of type Ia, ∼ 1 week before maximum. Strong lines of Si II, S II, Fe III, and Ca II are present, with minima measured at 618.2, 585.0, 549.8, 534.0, 492.3, 434.0, 403.8, and 372.2 nm. The photospheric expansion velocity, measured from the minimum of the Si II doublet, is ∼ 13 300 km/s, while that measured from Ca II is ∼ 22 000 km/s, making this object more similar to SNe 1992A and 1981B than to SN 1994D (Patat et al. 1996, 278, 111). Strong interstellar Na I D absorptions are measured at 599.2 and 599.9 nm — i.e., at the same recession velocity (∼ 5250 km/s) of the narrow Hα, [N II] and [S II] emissions of the parent galaxy. On Oct. 24.14, the following magnitudes were measured with the Dutch 0.9-m telescope: B = 16.32, B − V = +0.32, V − R = +0.27; and the position of SN 1996bo was measured to be 5′′.3 east and 1′′.7 south of the galaxy’s center.”

COMET 126P/IRAS Total visual magnitude estimates: Oct. 8.22 UT, 12.4 (R. Keen, Mt.

Thorodin, CO, 0.32-m reflector); 11.91, 11.7 (J. Carvajal, Madrid, Spain, 0.21-m reflector); 13.78, 12.3 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); 15.47, 11.8 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector).

1996 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green